using elimination method how do i solve this
0.05x+0.05y=44
0.15x+0.05y=48

Answer by amoresroy(361)

using elimination method how do i solve this
0.05x+0.05y=44
0.15x+0.05y=48
Dubtract 2nd equation by 1st equation to eliminate y
.15x-.05x+.05y-.05y = 48-44
.10x = 4
x = 40
Substitute x = 40 in 1st equation
.05(40) + .05y = 44
.05y = 44-2
y = 840